lband wide calibration


     The lband wide receiver covers 1120 to 1730 Mhz. It is a native linear receiver with an optional hybrid than can form circular polarizations (located outside of the dewar).

recent system performance measurements
Daily monitoring of Tsys
Dewar temperatures
gain curves
lband wide measurements by date
cal values
rfi measurements
lband wide measurements by topic:
Resonances Cal jumps Rf filterbank The hybrid
Compression  dewar Warmup

plot/images entire


History  (top)

Gain curves: (top)

180625: post hurricane maria gain curves.. dec17 - jun18.
02nov03: lbw gain curves from mar03 thru oct03  (gain curves)
17feb03: new lband wide receiver installed.
lband wide gain curves after horn focus moved (data 21nov02 thru 31dec02). (gain curves)
gain, tsys calibration lbw (sep01). (gain curves.)
fit to gain(az,za) and comparison of pitch, roll, focus loss to measured losses (aug00). (gain curves.)

 Lband wide measurements by date:  (top)


190423: switching the polA,B  drive cables for the hp1370 filter caused the problem to go  away.
190329: lbw 1370hp polB filter always inserted in RF.
190207: PolA,B phase difference vs frequency using the correlated cal and stokes measurements.
190110: lbw 1370 hipass filter shows unstable power output when switched.


180708: post hurricane maria sefd curves.. dec17-jun18 (sefd curves)
180625: post hurricane maria gain curves.. dec17 - jun18. (gain curves)


05may17: polA jumps (about 2% of Tsys)
28mar17: a3160 has gain variations in polB. rf filter bank switches failing.


21jul16: new cal values installed that were measured back on 28apr16.
29jun16: SEFD changes jan2006 through jun2016
22apr16: sefd vs frequency for data jan14 -> apr16
22apr16: calibration run after painting shutdown to check pointing, system performance


22sep15: lbw PolB pwr changes rapidly (35%) in 50 secs as the dewar 16K stage changes by 8 K
18sep15: lbw polB stable after working on dewar
31aug15: lbw polB drifting when dewar warmed up a little
19aug15: polB drifting around.. a2812 reports problems..
13jun15: calibratio runs after lbw reinstalled in dome
08apr15: a2703, polB jumps (cable connectors)


21apr14: a2703,a2808 report bad baselines...
22apr13: check filter bank by switching away and then back to filter.
08apr13: a2749 had polB drift by 2% for the first 45 minutes of run.
13jan13: calibration results: jun12-dec12. updated pointing offset.


12dec12: oct-nov12 Tsys increase. calibration data shows it's not a problem with the cals.
01dec12: 16k lbw cooldown took > 36 hours
01nov12: polA ripple from 1370 hipass rf switches.
12aug12 Cal values remeasured, installed. Data shows that calPolA jumped on 07feb12.


15aug11: 1360 narrow band rfi
22mar11 - a2335.. total power plots for some of the mar11 data. (total power drifts during on,offs)
14mar11- 1357.2 Mhz rfi.
30-31jan11: polB baseline ripple seen by a2335


28oct10: Tsys jumps after warmup,cooldown sequence.
11oct10: monitor lbw stability with 3 1800 sec drift scans.
07oct10: a2335 sees instabilities in polA and polB.
29sep10: 4200 sec integration to check stability.
28sep10: 1340 Mhz tone coming from radar blanker lo.
27sep10: scan lbw band.. images, avg spec,rms/mean, and tp vs time.
09sep10: instabilities in polB seen by a2335 in jul10.
16mar10: crosshead rfi in the lbw feed (1.2 hz periodicity).
08feb10:3C48 in Degrees K across entire band.


08nov09: 1375 mhz rfi.
08may09: lbw dewar warmup time.
05may09: rfi from lbw crosshead.


27nov08: trying to remove the lbw 1400 Mhz resonance in software
26sep08: lbw cals remeasured.
29aug08:bw cals remeasured before warmup
02jul08: lbw cooldowns, omt stage not getting below 180K
22apr08: pol B is noisier than polA.
08apr08: lbw cal box brought to lab and components swept on network analyzer (.pdf)
02apr08: cal values measured on 21apr07 finally installed.


18dec07: sefd shows polB Tsys 25%  greater than polA.
21apr07: polA resonance at 1400 Mhz gone after omt bolts tightened. PolB at 1395 still present.
19feb07: lbw beam  maps of 3C138.
13feb07: sefd around 1100 Mhz and some rfi plots.


30oct06 cable test lbw in upstairs iflo shows loose cable in 750 mixer chassis.
09mar06: cycle through the 9 rf filters checking Tsys variations
07mar06: lbw calPolB. tapping the cables external to dewar to not change caldeflection.
07mar06: 1230-1470Mhz filter. See if Tsys stable when in, out.
07mar06: jun05 warmup/cooldown caused resonance in polA and polB.


26dec05: no power in  polB for 2 hours (probably and iflo problem).04dec05: lbw cal value polB jumps by 20%.
28nov05: cal values for new diodes installed.
26nov05: lbw polB filter 3  shows  jumps (1230-1470 Mhz)
28oct05: lbw calibration jumps for murray's OH data.
05sep05: lbw receiver status jun05-sep05. 1400Mhz resonance and diode 1 dies.
23jun05: resonance in lbw polA at 1400 Mhz.
21jun05: installed new cal values after polB cal cable fixed.
03jun05: time for lbw dewar to cool down.
16may05: power jumps, 1.2 hz gain variation from crosshead. Something loose in dewar.


02aug04: ripples in lbw  around 1390,1400 Mhz. seen a1861(may04), a1914(aug04)
23apr04: polB cal jumps.
09apr04: system performance after kevlar cables tensioned and tertiary moved.

2003 Images/plots of the entire band

02nov03: lbw gain curves from mar03 thru oct03  (gain curves)
11jul03: compare lbw horns: kildal and german cortes.
9apr03:IFLO compression by the FAA 1350/1330 radars using lbw,lbn (29apr03).
17apr03:  LBW SEFD vs frequency using 3C138.
17apr03: resonances in the new lbw receiver.
03mar03 first look newlbw performance.
Tsys, Trcvr, and scattered radiation from hilltop measurements.
17feb03 new lband wide receiver installed


lband wide gain curves after horn focus moved (data 21nov02 thru 31dec02). (gain curves)
lbw system performance after horn focus move (21nov02 thru 31dec02).
lbw gain increase after horn moved to correct phase center (22nov02)
100 sec oscillation in polB total power (2001,2002)
Compare sefd 01sep01 thru 01may02 lbw,lbn using new cal values (may02)
resonances in the lband wide OMT (mar02)


gain, tsys calibration lbw (sep01). (gain curves.)
problems using the lbw system in circular mode (hybrid in) jul01.


fit to gain(az,za) and comparison of pitch, roll, focus loss to measured losses (aug00). (gain curves.)
lband Sefd, tsys, and gain (night) aug00 (html)
    (position switching on flux calibrators using the correlator).
Compare Sefd, gain, Tsys at multiple lbw frequencies (14jul00 daytime)
Compare lbw and lbn receivers (Sefd, gain, and tsys  12jul00 daytime).

Lbw measurements by topic: (top)


            Images/plots of the entire band


Cal jumps:

Rf filterbank

The hybrid:


Warmup/cooldown times for dewar: